OpenStudy (anonymous):

Help figuring out how to do the equation 5c - c + 10 = 34

OpenStudy (princeharryyy):

5c-c = 4c => equation can be written as 4c + 10 = 34 now take 10 to the right side its sign will get changed equation then can be written as ..... 4c = 34 - 10 =>4c = 24 now divide 24/4 = 6 => c= 6 @yololifelove

OpenStudy (anonymous):

yay thanks :) can I ask one more 4(x + 1) + 8 = 24

OpenStudy (princeharryyy):

ohk

OpenStudy (princeharryyy):

4x + 4 +8 =24 => 4x + 12 = 24 =>4x = 24-12 => 4x =12 => x =3

OpenStudy (princeharryyy):

@yololifelove

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

thanks :)

OpenStudy (princeharryyy):

welcome ))) @yololifelove